March 11th 2022

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • two VTP installed in crates, will need to be cabled waiting for power outage, might be outage on Monday
    • CDet : taking cosmics for left side
      • VETROC
      • ADC for CDet
      • SIDIS
    • GEn RP
  • Hodoscope : mostly focusing on analysis, planning for travel, plan for work on HV, spare crate shipped from Glasgow
  • RICH : box open yesterday, still working on detector. No DAQ required for now. VETROC , use VXS crate + VETROC , need to coordinate with hodoscope, GRINCH, CDET for NINO - 2000 channels
  • FADC : could use HRS FADC
    • 14 FADCs
  • SBS GEM : in May June
  • HV : need to get new RPi to Bryan